OpenStudy (anonymous):

If f(t)=t^3 +2, g(x)=2x+3, p(x)=(f(g(x)), and r(t)=f(g(t)). Compute p(-1) and r(-2).

OpenStudy (unklerhaukus):

to get p(-1), we first have to work out p(x), p(x)=f(g(x)) what do you get when you substitute the definition of g(x) into this

OpenStudy (unklerhaukus):

p(x)=f( )

OpenStudy (anonymous):

Okay I'm not sure what you mean, still new to functions

OpenStudy (unklerhaukus):

yeah i understand , the notation is a bit different to stuff you have learnt before, ill give you a different example, and hopefully you can see what im doing

OpenStudy (unklerhaukus):

say h(x)= 100x , this is exactly the same as h(y)=100y and also exactly the same as h(z)=100z the function h is just multiplying what ever is in the brackets by 100, so you could even say that h(cat)= 100cat

OpenStudy (anonymous):

Okay, I follow

OpenStudy (unklerhaukus):

now say that j(x)= 55+ x [same thing as j(cat)= 55 +cat ]

OpenStudy (unklerhaukus):

now we can put a more complicated thing in the bracket of the function lets put the function h as the input for j j(h(x))=j(100x) now applying the definition of the j function , we get j(100x)= 55+(100x)

OpenStudy (unklerhaukus):

or we could even put the function j as the input to the function j j(j(t))= 55 + j(t) = 55 + (55+t) = 110 + t am i making sense?

OpenStudy (anonymous):

Yes, so far so good

OpenStudy (unklerhaukus):

cool, now we can even use numbers as inputs j(j(9))= 55+j(9) = 55+55+9 =119

OpenStudy (unklerhaukus):

______________ ______________ now if you can follow all this you should be able to express p(x) [p as a function of x] f(t) = t^3 +2, g(x) = 2x+3, p(x) = f(g(x)) = f( ) =

OpenStudy (anonymous):

f(2x+3)

OpenStudy (unklerhaukus):

good, now apply the definition of the f function (cube the input, and add two)

OpenStudy (anonymous):

8+5?

OpenStudy (unklerhaukus):

should still have x's in it

OpenStudy (unklerhaukus):

[note that you dont have to simplify the cubed term leave it in brackets]

OpenStudy (anonymous):

2x +5 (t^3)?

OpenStudy (unklerhaukus):

f(t) = t^3 +2, g(x) = 2x+3, p(x) = f(g(x)) = f(2x+3) = ( )^3+2

OpenStudy (anonymous):

p(x)=(2x+3)^3 +2

OpenStudy (unklerhaukus):

right, so you have an expression for p as a function x now we can compute p(-1) just put the -1 as the input for this definition of p and then we can simplify the numbers

OpenStudy (anonymous):

P(-1)= (-2 + -3)^3 +2 =-123

OpenStudy (unklerhaukus):

that would right [if p(x)=(2x-3)^3 +2] but its a +3 [not -3] p(x)=(2x+3)^3 +2

OpenStudy (anonymous):

P(-1) =

OpenStudy (anonymous):

oops 1 second

OpenStudy (anonymous):

P(-1) = 3

OpenStudy (unklerhaukus):

YES!! very good √

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

Is the r similar?

OpenStudy (unklerhaukus):

For r(-2). yes its the same idea , f(t) = t^3 +2 g(x) = 2x+3, r(t) = f(g(t)) = = r(-2) =

OpenStudy (anonymous):

is g(x) equivalent to g(t)?

OpenStudy (anonymous):

r(-2)=(2x-3)^3 +2?

OpenStudy (unklerhaukus):

g(x) is equivalent to g(t)

OpenStudy (anonymous):

so then f(2x+3) = (2x+3)^3 +2

OpenStudy (unklerhaukus):

yes actually you can see that p and r are the same function

OpenStudy (anonymous):

Oh okay

OpenStudy (unklerhaukus):

so you just need to get r(-2)=p(-2) and you already worked out p(x)

OpenStudy (anonymous):

so r(-2) = 1

OpenStudy (kainui):

The important thing to take away from this is that any time you have a function, we just call it a random letter like f or g or whatever and then in parenthesis you put another letter showing what you can replace in that function. f(x) just means the function's name is f and anywhere you see an x in the function, you replace it. You can replace the x with anything. A number. Another letter. A whole other equation. Or maybe an elephant. So if you have f(x)=x and you want to replace x with 2, then you have f(2)=2. Just replacing stuff. If you want to replace x with (x^2-1) then f(x^2-1)=(x^2-1). Another example, suppose you have g(t)=t^2+5t and you want to replace t with elephants? No problem, g(elephants)=(elephants)^2+5(elephants).

OpenStudy (anonymous):

Thanks that makes sense

OpenStudy (unklerhaukus):

+3 [not -3] sorry, i have to go catch a train if you still can't get this one , you might like to close this question and open a new one, just for r(-2) and if you show your working somelse will be able to help

OpenStudy (anonymous):

Okay, thanks for all your help